Question 11

Using the financial information below, calculate company XYZ's ROA (adjusted for debt financing). Enter your answer as a decimal with four places of precision (i.e. 0.1234).

xid-100948877_1
Answer #1

Formula to calculate Return On Asset, ROA = Net Income/Average Assets

where, Net Income = net earnings/net income in a financial year

   Average Assets = present year's assets plus previous year's assets divided by 2

Given, Net Income = 2000

  Average Assets = (32000+26500)/2 = 29250

ROA = 2000/29250 = 0.0684 = 6.84%

Hence the answer is = 0.0684

This means that the profit earned per rupee by the company is 6.84 paise.